# SkillMimic: Learning Reusable Basketball Skills from Demonstrations

### v1

@@ 5.2. Learning Skills by Imitation
The state observed by the skill policy is not under direct supervision and may theoretically encompass any information that is available in the simulation environment. Importantly, this information does not necessarily have to be included in the reference HOI data. Similar to prior arts (Peng et al., 2022), we transform all coordinates into the root local coordinate of the humanoid, which aligns the data distribution and benefits the generalization performance. For the humanoid, we observe its global root height, local body position, rotation, position velocity, and angular velocity. These representations form the humanoid proprioception $\boldsymbol{o}^{prop}_{t}$. In addition, we detect net contact forces $\boldsymbol{o}^{f}_{t}$ for all fingertips, which helps to sense contact and accelerate training. For the object, we observe its local position, rotation, velocity, and angular velocity, which form the object observation $\boldsymbol{o}^{obj}_{t}$. Finally, the state perceived by the skill policy is
(1) $\boldsymbol{s}_{t}=\{\boldsymbol{o}^{prop}_{t},\boldsymbol{o}^{f}_{t},\boldsymbol{o}^{obj}_{t}\}.$
@@
Skill labels $\boldsymbol{c}_{j}$ are also used as inputs to the policy to differentiate between various skills, which is crucial for skill reuse and switching. We use one-hot encodings to represent the skill label.
5.2.3. Policy and Action.
@@
The policy output is modeled as a Gaussian distribution with dimensions equal to the DOF number of the humanoid robot, featuring constant variance. The mean is modeled by a three-layer MLP consisting of [1024, 512, 512] units with ReLU activations. We use the action $\boldsymbol{a}_{t}$ sampled from the policy as the target joint rotations for a full set of PD controllers. The PD controllers adjust and output the joint torques to reach the target rotations.
5.2.4. Physical Simulation of the Humanoid, Ball, and Environment.
We use Isaac Gym (Makoviychuk et al., 2021) as the physics simulation platform. For GRAB (Taheri et al., 2020) and BallPlay-V, the whole-body humanoid follows the SMPL-X (Pavlakos et al., 2019) kinematic tree and has a total of 52 body parts and 51$\times$3 DOF actuators where 30$\times$3 DOF is for the hands and 21$\times$3 DOF for the rest of the body. For BallPlay-M, the humanoid model consists of 53 body parts and 52$\times$3 DOF actuators, the hands having 30$\times$3 DOF and the rest of the body having 22$\times$3 DOF. The basketball is modeled as a sphere with a radius of 12 cm, which is close to the size of a real-world basketball. The restitution coefficients for the plane and the ball are set to 0.8 and 0.81, respectively, to ensure the ball’s bounce closely resembles real-world basketball behavior. The humanoid’s mass is set to match that of a real player. We set the ball’s density to 1000 kg/m3 to enhance stability and accelerate training convergence, while other physical parameters remain at their default settings. Despite being trained with fixed physical properties, our method can withstand a wide range of changes in physical properties during inference, such as variations in the ball’s density, radius, and restitution, as shown in Tab. 2.
5.3. Unified HOI Imitation Reward
The design of the unified HOI imitation reward in Fig. 3 (b) is crucial to the performance of learning interaction skills. This reward function is pivotal as it must not only accurately gauge the efficacy of skill acquisition but also avoid the incorporation of skill-specific settings to maintain robust cross-skill generalizability.
Traditional imitation rewards often focus solely on imitating human motions (Peng et al., 2018, 2021, 2022) or neglect interactive contacts (Zhang et al., 2023a), leading to suboptimal imitation performance and necessitating individualized hyperparameter tuning. Recognizing the interlocking relationship between interaction and contact, we propose a straightforward and universal contact modeling approach called Contact Graph. A corresponding CG imitation reward $r_{t}^{cg}$ is designed to enhance the precision of interaction imitation. Beyond the standard body motion imitation reward $r_{t}^{b}$ and object motion imitation reward $r_{t}^{o}$, we incorporate a relative position reward $r_{t}^{rel}$ to further improve interaction accuracy. Additionally, we propose an adaptive velocity regularization term $r_{t}^{reg}$ to suppress high-frequency jitters. These sub-rewards are multiplied to encourage balanced learning, thereby avoiding local optima, as justified in Tab. 1. The combination of these innovations forms a unified reward function, enabling the learning of diverse basketball skills in BallPlay-M with the same configuration. To conclude, the complete HOI imitation reward is
(2) $r_{t}=r_{t}^{b}*r_{t}^{o}*r_{t}^{rel}*r_{t}^{reg}*r_{t}^{cg}.$
Next, we will explore the methodology involved in crafting these reward functions.
5.3.1. Contact Graph.
We observe that existing imitation rewards (Peng et al., 2021, 2022; Zhang et al., 2023a) only measure kinematic properties and are insufficient to measure precise contacts. To tackle this problem, we propose the Contact Graph (CG) to model the contact in general interactions and design a CG imitation reward.
As illustrated in Fig. 4, the CG is a complete graph where every pair of distinct nodes is connected by a unique edge, defined as $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}\in\{0,1\}^{k}$ is the set of $k$ nodes and $\mathcal{E}\in\{0,1\}^{k(k-1)/2}$ is the set of edges. A CG node stores a binary value indicating whether it contacts other nodes. Each CG edge stores a binary label that denotes the contact between two nodes, where 1 represents contact, and 0 means no contact. CG node and edge values are calculated frame by frame, explicitly describing the mutual contact relationship between CG nodes at different moments. Contacts inside a node are not considered. In practice, we can use the nodes set $\mathcal{V}$ or edge set $\mathcal{E}$ depending on the needs.
The definitions of nodes are flexible. For example, a node can be a single part (e.g., a fingertip), or an aggregation of multiple parts (e.g., the whole left hand). Hinged objects can also be broken up into finer nodes (Fan et al., 2023; Geng et al., 2023; Zhang et al., 2024). The definition of CG is unified for a certain scene and shared between diverse interaction skills. For example, in basketball scenarios, we can aggregate two hands as a node, the rest of the bodies as a node, and the ball as a node, as illustrated in Fig. 4. This simple CG is effective enough for learning all the basketball skills covered in BallPlay-M. Fig. 6 shows a subset of these skills.
Theoretically, the reference CG edge values can be extracted from accurate kinematic HOI data by calculating the mesh collision or distances. A CG node value can be yielded by the logic ”OR” operation on all its connected edges.
@@ 5.3. Unified HOI Imitation Reward
Figure 4. We propose the Contact Graph (CG) to model general contacts within an explicitly defined scene. The node stores a binary value that denotes whether it contacts other nodes. Each edge stores a binary value indicating whether the two connected nodes are in contact. The node definition is unified for a certain scene and shared between diverse interactive skills. For example, we define three nodes: hands, hands-exclusive body, and ball, to form a simple CG to model contacts for diverse basketball skills.
@@
Figure 5. The HOI imitation falls into kinematic local-optimal solutions without Contact Graph Reward (CGR): (b) use the head to help control the ball; (e) use the wrist to contact the ball; (h) fail to catch the object; (k) support the table to keep balance. In comparison, the guidance of CGR effectively yields precise interactions, as shown in (c, f, i, l).
5.3.2. Contact Graph Reward.
Kinematic imitation rewards (Peng et al., 2018, 2021; Zhang et al., 2023a) fall short in measuring contacts and thus yield poor performance in HOI imitation. For example, during the toss skill training (Fig. 5 (a)), the contact between the right hand and the ball will in most cases cause the ball to drop, making the kinematic rewards drastically smaller. In this case, the policy may learn a local-optimal solution that uses the head and left hands to stabilize the ball, as demonstrated in Fig. 5 (b). To tackle this problem, we propose the Contact Graph Reward (CGR) as a critical complement of kinematic imitation rewards to learn precise contact imitation. The CG error is defined as $\boldsymbol{e}_{t}^{cg}=|\boldsymbol{s}_{t}^{cg}-\hat{\boldsymbol{s}}_{t}^{cg}|$, where ||$\cdot$|| ⋅ | calculates element-wise absolute value, $\boldsymbol{s}^{cg}_{t}$ and $\hat{\boldsymbol{s}}^{cg}_{t}$ denotes the simulated and reference CG state respectively, which can be the nodes set $\mathcal{V}_{t}\in\{0,1\}^{k}$ or the edge set $\mathcal{E}_{t}\in\{0,1\}^{k(k-1)/2}$ depending on the needs. The CGR is measured by CG error, with independent weights on different elements:
(3) $r_{t}^{cg}=\text{exp}(-\sum_{j=1}^{J}\boldsymbol{\lambda}^{cg}[j]*\boldsymbol{e}_{t}^{cg}[j]),$
where $\boldsymbol{e}_{t}^{cg}[j]$ is the $j$th element of $\boldsymbol{e}_{t}^{cg}\in\{0,1\}^{J}$, a binary label representing a contact error; $\boldsymbol{\lambda}^{cg}[j]$ is the $j$th element of $\boldsymbol{\lambda}^{cg}\in\mathbb{R}^{J}$, a hyperparameter controls the sensitivity of a contact. Fig. 5 and Tab. 1 justify the effectiveness of CGR, where experiments without CGR show kinematic local-optimal solutions while using CGR effectively eliminates such problems.
5.3.3. Kinematic Rewards.
Kinematic imitation rewards form the core of the HOI imitation. We design these rewards in four distinct parts: the Body Kinematics Reward $r_{t}^{b}$, the Object Kinematics Reward $r_{t}^{o}$, the Relative Motion Reward $r_{t}^{rel}$, and a Velocity Regularization term $r_{t}^{reg}$.
(4) $r_{t}^{b}=r_{t}^{p}*r_{t}^{r}*r_{t}^{pv}*r_{t}^{rv},$
(5) $r_{t}^{p}=\text{exp}(-\lambda^{p}*e_{t}^{p}),\quad e_{t}^{p}=\text{MSE}(\boldsymbol{s}_{t}^{p},\hat{\boldsymbol{s}}_{t}^{p}),$
(6) $r_{t}^{o}=r_{t}^{op}*r_{t}^{or}*r_{t}^{opv}*r_{t}^{orv},$
The relative motion is represented as a vector group, obtained by subtracting the object’s position from the key body positions. The Relative Motion Reward $r_{t}^{rel}$ is also calculated following Eq. 5 and is effective in constraining the relative motion between the object and key body points to be consistent with the reference.
(7) $r_{t}^{reg}=\text{exp}(-\lambda^{reg}*e_{t}^{acc}),\quad e_{t}^{acc}=\text{mean}\left(\frac{||\boldsymbol{s}_{t}^{acc}||^{2}}{||\hat{\boldsymbol{s}}_{t}^{vel}||^{2}+\lambda^{reg}}\right),$
5.4. Reusing Acquired Skills for Complex Tasks
Although the definition of skills can be quite broad, employing more complex data to define more intricate skills introduces additional demands for data collection. Instead, we consider leveraging the skills acquired through SkillMimic to accomplish long-term, complex tasks. Specifically, we propose training a High-Level Controller (HLC) to reuse acquired skills for implementing complex tasks. As depicted in Fig. 3 (c), the HLC takes as input the current HOI state $\boldsymbol{s}_{t}$ and task observation $\boldsymbol{h}_{t}$, and outputs a discrete skill embedding $\boldsymbol{c}_{t}$ that serves as inputs to a pre-trained SkillMimic policy, which subsequently generates actions for the humanoid. Compared to previous hierarchical approaches that train a High-Level Controller (HLC) to reuse motion priors (Peng et al., 2022; Zhang et al., 2023b), our method does not require rewards designed for interactions. Instead, simple goal-related rewards can be highly effective. This significantly simplifies the design of task rewards and greatly accelerates the convergence speed of training. In the following sections, we present four representative complex tasks and our detailed solutions. Results in Tab. 4 and Fig. 11 justify the superiority of our method.
5.4.1. Throwing
(8) $r_{t}^{throwing}=\text{exp}(-|h_{t}^{ball}-2.5|),$
where $h_{t}^{ball}$ is the ball height.
5.4.2. Heading
(9) $r_{t}^{heading}=\text{exp}(-||\boldsymbol{x}_{t}^{ball}-\boldsymbol{x}_{t}^{target}||^{2}),$
5.4.3. Circling
(10) $r_{t}^{circling}=r_{t}^{v}*\text{exp}(-|d^{target}-||\boldsymbol{x}_{t}^{ball}-\boldsymbol{x}_{t}^{center}||^{2}|),$
(11) $r_{t}^{v}=\begin{cases}1,&\text{if}\quad||\boldsymbol{v}_{t}^{ball}||^{2}>0.5\\ 0.1,&\text{else}\\ \end{cases}$
where $\boldsymbol{v}_{t}^{ball}$ is the ball velocity.
5.4.4. Scoring
(12) $r_{t}^{scoring}=r_{t}^{v}*(r_{t}^{heading}+r_{t}^{bonus}+0.2*r_{t}^{throwing}),$
(13) $r_{t}^{bonus}=\begin{cases}1,&\text{if scored}\\ 0,&\text{else}\\ \end{cases}$
After only three hours of training (on a Nvidia RTX 4090 GPU), the HLC is capable of controlling a humanoid to dribble the ball towards the basket, execute layups at the right moments to score, retrieve the rebound, dribble away, and repeat the scoring sequence.
(a) Jump Shot
(b) Rebound
(c) Layup
(d) Turnaround Layup
(e) Dribble Left
(f) Dribble Right
(g) Dribble Forward
(h) Dribble Turn
(i) Pass
(j) Pickup
(k) Cross Leg
(l) Back Dribble
@@ 5.4.4. Scoring
Figure 6. Simulated humanoids exhibit comprehensive basketball skills. SkillMimic can teach humanoids a wide range of basketball skills using the same configuration in a purely data-driven manner, covering almost all fundamental basketball skills. Keyframes are placed in chronological order from left to right.
6. Experiment
Our experiments are divided into two main parts. The first part evaluates skill learning (Sec. 6.3). This includes comparisons with existing methods, robustness testing of the acquired skills, and ablation studies on the reward function designs and data volume. The second part assesses learning complex tasks by reusing acquired skills (Sec. 6.4), including comparisons with existing methods and a demo of a two-player competition.
6.1. Experimental Setup
We use Isaac Gym (Makoviychuk et al., 2021) as the physics simulation platform. All experiments are trained on a single Nvidia RTX 3090 or 4090 GPU, with 2048 parallel environments. For GRAB and BallPlay-V, both the simulation and PD controller operate at 60 Hz, while the skill policy is sampled at 30 Hz. For BallPlay-M, the simulation and PD controller run at 120 Hz, with the skill policy sampled at 60 Hz. We resample the reference HOI clips to match the skill policy frequency, and the high-level policy is sampled at 20 fps. All neural networks are implemented using PyTorch and trained using Proximal Policy Optimization (Schulman et al., 2017). We use the edge set $\mathcal{E}$ of the CG and calculate the CG edge values by judging the contact force of each CG node. The setting of hyperparameters is fixed for all experiments and can be found in the appendix.
Table 1. Ablation study on imitation reward design. We experiment on two HOI datasets, including GRAB, which contains multiple objects, and BallPlay-V, which has significant data inaccuracies. The result highlights the importance of the CG reward and multiplication in achieving precise HOI imitation.

### v2

@@ 3.3.2 HOI Observation
The state observed by the IS policy is not under direct supervision and may theoretically encompass any information that is available in the simulation environment. Importantly, this information does not necessarily have to be included in the reference HOI data. Similar to prior arts [38], we transform all coordinates into the root local coordinate of the humanoid, which aligns the data distribution and benefits the generalization performance. For the humanoid, we observe its global root height, local body position, rotation, position velocity, and angular velocity. These representations form the humanoid proprioception $\boldsymbol{o}^{prop}_{t}$. In addition, we detect net contact forces $\boldsymbol{o}^{f}_{t}$ for all fingertips, which helps to sense contact and accelerate training. For the object, we observe its local position, rotation, velocity, and angular velocity, which form the object observation $\boldsymbol{o}^{obj}_{t}$. Finally, the state perceived by the IS policy is
$\boldsymbol{s}_{t}=\{\boldsymbol{o}^{prop}_{t},\boldsymbol{o}^{f}_{t},\boldsymbol{o}^{obj}_{t}\}.$ (1)
@@
Skill labels $\boldsymbol{c}_{j}$ are also used as inputs to the IS policy to differentiate different skills. We use one-hot encodings to represent the skill label.
3.3.3 IS Policy and Action
@@
The policy output is modeled as a Gaussian distribution with dimensions equal to the DOF number of the humanoid robot, featuring constant variance. The mean is modeled by a three-layer MLP consisting of [1024, 512, 512] units with ReLU [33] activations. We use the action $\boldsymbol{a}_{t}$ sampled from the policy as the target joint rotations for a full set of PD controllers. The PD controllers adjust and output the joint torques to reach the target rotations.
@@
Figure 4: We propose the Contact Graph (CG) to model general contacts within an explicitly defined scene. The node stores a binary value that denotes whether it contacts other nodes. Each edge stores a binary value indicating whether the two connected nodes are in contact. The node definition is unified for a certain scene and shared between diverse interactive skills. For example, we define three nodes: hands, hands-exclusive body, and ball, to form a simple CG to model contacts for diverse basketball skills.
3.4 Unified HOI Imitation Reward
This HOI imitation reward function is pivotal as it must not only accurately gauge the efficacy of interaction skill acquisition but also avoid the incorporation of skill-specific settings to maintain robust cross-skill generalizability.
Classic imitation rewards often focus solely on imitating locomotions [36], lacking an established baseline for HOI imitation. Although adversarial imitation rewards [37, 38] are effective in locomotion tasks, we find that they perform poorly in HOI imitation. This may be because adversarial imitation rewards are relatively coarse-grained, making it difficult for them to provide the precise guidance needed for accurate interaction learning.
The body kinematics imitation reward $r_{t}^{b}$ encourages the alignment of body movements by considering joint position, rotation, and their velocities. Similarly, the object kinematics imitation reward $r_{t}^{o}$ ensures consistent object movements with the reference. However, solely relying on these two rewards often leads to unstable learning and unnatural motions. We incorporate a relative motion reward $r_{t}^{rel}$ to highlight the relative relationships in interactions, which is crucial for interaction naturalness. Recognizing the interlocking relationship between interaction and contact, we propose a straightforward and universal contact modeling approach called Contact Graph (CG). A corresponding CG imitation reward $r_{t}^{cg}$ is designed to enhance the precision of interaction imitation. Additionally, we propose an adaptive velocity regularization term $r_{t}^{reg}$ to suppress high-frequency jitters. These sub-rewards are multiplied rather than added to encourage balanced learning, thereby avoiding local optima, as justified in Tab. 1. The combination of these innovations forms a unified imitation reward, enabling the learning of diverse basketball skills in BallPlay-M with the same configuration. To conclude, the complete HOI imitation reward is
$r_{t}=r_{t}^{b}*r_{t}^{o}*r_{t}^{rel}*r_{t}^{reg}*r_{t}^{cg}.$ (2)
Next, we will discuss the methodology involved in crafting the CG reward $r_{t}^{cg}$. Details about the kinematic imitation reward $r_{t}^{b}*r_{t}^{o}*r_{t}^{rel}*r_{t}^{reg}$ are in the appendix.
3.4.1 Contact Graph
We observe that existing imitation rewards [37, 38, 57] only measure kinematic properties and are insufficient to measure precise contacts. To tackle this problem, we propose the Contact Graph (CG) to model the contact in general interactions and design a CG imitation reward.
As illustrated in Fig. 4, the CG is a complete graph where every pair of distinct nodes is connected by a unique edge, defined as $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}\in\{0,1\}^{k}$ is the set of $k$ nodes and $\mathcal{E}\in\{0,1\}^{k(k-1)/2}$ is the set of edges. A CG node stores a binary value indicating whether it contacts other nodes. Each CG edge stores a binary label that denotes the contact between two nodes, where 1 represents contact, and 0 means no contact. CG node and edge values are calculated frame by frame, explicitly describing the mutual contact relationship between CG nodes at different moments. Contacts inside a node are not considered. We use the edge set $\mathcal{E}$ for imitation learning.
The definitions of nodes are flexible. For example, a node can be a single part (e.g., a fingertip), or an aggregation of multiple parts (e.g., the whole left hand). Hinged objects can also be broken up into finer nodes [10, 14, 56]. The definition of CG is unified for a certain scene and shared between diverse interaction skills. For example, in basketball scenarios, we can aggregate two hands as a node, the rest of the bodies as a node, and the ball as a node, as illustrated in Fig. 4. This simple CG is effective enough for learning all the basketball skills covered in BallPlay-M.
@@ 3.4.1 Contact Graph
Figure 5: The HOI imitation falls into kinematic local-optimal solutions without Contact Graph Reward (CGR): (b) use the head to help control the ball; (e) use the wrist to contact the ball; (h) fail to catch the object; (k) support the table to keep balance. In comparison, the guidance of CGR effectively yields precise interactions, as shown in (c, f, i, l).
(a) Layup
(b) Turnaround Layup
(c) Dribble Right
(d) Dribble Forward
(e) Dribble Turn
(f) Jump Shot
(g) Rebound
(h) Pass
(i) Pickup
(j) Cross Leg
@@
Figure 6: Simulated humanoids exhibit comprehensive basketball skills. SkillMimic can teach humanoids a wide range of basketball skills using the same configuration in a purely data-driven manner, covering almost all fundamental basketball skills. Keyframes are placed in chronological order from left to right.
3.4.2 Contact Graph Reward
Kinematic imitation rewards [36, 37, 57] fall short in measuring contacts and thus yield poor performance in HOI imitation. For example, during the toss skill training (Fig. 5 (a)), the contact between the right hand and the ball usually cause the ball to drop, making the kinematic rewards drastically smaller. In this case, the policy may learn a local-optimal solution that uses the head and left hands to stabilize the ball, as demonstrated in Fig. 5 (b). To tackle this problem, we propose the Contact Graph Reward (CGR) as a critical complement of kinematic imitation rewards to learn precise contact imitation. The CG error is defined as $\boldsymbol{e}_{t}^{cg}=|\boldsymbol{s}_{t}^{cg}-\hat{\boldsymbol{s}}_{t}^{cg}|$, where ||$\cdot$|| ⋅ | calculates element-wise absolute value, $\boldsymbol{s}^{cg}_{t}$ and $\hat{\boldsymbol{s}}^{cg}_{t}$ denotes the simulated and reference CG state respectively, which is the edge set $\mathcal{E}_{t}\in\{0,1\}^{k(k-1)/2}$. The CGR is measured by CG error, with independent weights on different elements:
$r_{t}^{cg}=\text{exp}(-\sum_{j=1}^{J}\boldsymbol{\lambda}^{cg}[j]*\boldsymbol{e}_{t}^{cg}[j]),$ (3)
where $\boldsymbol{e}_{t}^{cg}[j]$ is the $j$th element of $\boldsymbol{e}_{t}^{cg}\in\{0,1\}^{J}$, representing a contact error; $\boldsymbol{\lambda}^{cg}[j]$ is the $j$th element of $\boldsymbol{\lambda}^{cg}\in\mathbb{R}^{J}$, a hyperparameter controls the sensitivity of a contact. Fig. 5 and Tab. 1 justify the effectiveness of CGR, where experiments without CGR show inaccurate interactions while using CGR effectively eliminates such problems.
3.5 Reusing Interaction Skills for High-level Tasks
In this section, we propose training a High-Level Controller (HLC) to reuse the interaction skills acquired through SkillMimic to accomplish high-level tasks. As depicted in Fig. 3 (c), the HLC takes as input the current HOI state $\boldsymbol{s}_{t}$ and task observation $\boldsymbol{h}_{t}$, and outputs a discrete skill embedding $\boldsymbol{c}_{t}$ that serves as inputs to a pre-trained SkillMimic policy, which subsequently generates actions for the humanoid.
Compared to previous hierarchical approaches that are based on locomotion priors [38, 55], our method leverages interaction skills as prior, which significantly simplifies the design of task rewards and greatly accelerates training. Moreover, unlike their mandatory hierarchical structure, our approach is motivated by data efficiency - tasks can be alternatively solved through direct HOI imitation when sufficient demonstrations are available, as shown in pickup experiments in Sec. 4.2, Fig. 7, and Fig. 8.
We present four representative high-level tasks:
Figure 7: Demonstration of the pickup skill learned from 40 HOI motion clips. Yellow denotes the initial frame. Left: The humanoid picks up a stationary ball effortlessly. Middle: The humanoid intercepts a ball with random velocity. Right: The humanoid adjusts after missing the ball initially (the frame in blue) and successfully retrieves it on the second attempt, showcasing the potential for learning robust and generalizable skills through extensive data collection.